# Degrees of Freedom for Radiating Systems

### v1

@@ Abstract
Electromagnetic degrees of freedom are instrumental in antenna design, wireless communications, imaging, and scattering. Larger number of degrees of freedom enhances control in antenna design, influencing radiation patterns and directivity, while in communication systems, it links to spatial channels for increased data rates and reliability, and resolution in imaging. The correlation between computed degrees of freedom and physical quantities is not fully understood, prompting a comparison between classical estimates, Weyl’s law, modal expansions, and optimization techniques. In this paper, it is shown that NDoF for arbitrary shaped radiating structures approaches the shadow area measured in squared wavelengths.
Index Terms:
Antenna theory, electromagnetic theory, radiation modes, computational electromagnetics, degrees of freedom, capacity, inverse source problems.
@@ I Introduction
Electromagnetic degrees of freedom (DoF) are significant across a wide range of electromagnetic (EM) applications, including wireless communication systems, antenna design, imaging, and measurement techniques [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The number of DoF (NDoF) emerges as a crucial factor in achieving desired performance levels. For instance, in wireless communication systems, NDoF correlates with the number of channels or spatial dimensions utilized for signal transmission, thereby providing higher data rates and improved reliability, especially in technologies like multiple-input multiple-output (MIMO) systems and intelligent surfaces [9, 7]. In antenna design, a higher NDoF relates to maximum directivity [6] and improved control over radiation patterns. In the realm of inverse problems, NDoF is intricately linked to the amount of independent data available and the achievable resolution.
@@
The relationship between the NDoF and the electrical size of an object have been explored for at least half-a-century investigating resolution in imaging systems [11] and antenna systems [12]. Small electrical sizes exhibit few DoF while large sizes display numerous DoF and hence e.g., potentially higher directivity and narrow beamwidth. Analytically solvable cases such as spherical regions with radius $a$ offer a simple estimate of $2(ka)^{2}$ DoFs [2] based on the first $L\approx ka$ spherical modes for wavenumber $k=2\pi/\lambda$ and wavelength $\lambda$. The number of propagating waves in hollow rectangular waveguides [13] are also solvable with asymptotically $2\pi A/\lambda^{2}$ DoFs for a cross sectional area $A$. These results are traced back to Weyl more than a century ago, who studied the eigenvalue distribution of the Laplace operator [14, 15]. Alternative estimates based on Nyquist ($\lambda/2$) sampling resulting in $8A/\lambda^{2}$ DoFs have also been extensively used [2, 16, 4, 5]. Moreover, communication systems have recently been investigated for different configurations [9, 7], with $2\pi A/\lambda^{2}$ DoFs for planar apertures.
In this paper, Weyl’s law [14, 15] and its connection to NDoF are first discussed. Weyl’s law show that the NDoF for arbitrary shaped bodies scaled with the surface area measured in squared wavelength in the same manner as for spherical regions. These NDoF can be interpreted as a communication system between the body and the region outside the body. For antenna systems it is more common to consider radiating systems, where the antenna radiates to the far-field. Here, it is shown that these radiating NDoF approaches the shadow area of the region measured in squared wavelengths. This NDoF is also two times the number of significant characteristic modes [10]. The result is also identical the NDoF from Weyl’s law for the special case of convex shapes but differ for general shapes.
@@ II Weyl’s law and propagating modes
NDoF is most straight forward determined for the number of propagating modes in a hollow waveguide. Canonical shapes such as rectangular or circular cross sections can be solved analytically [13]. These and arbitrary shaped cross sections can also be treated as a special case of Weyl’s law [14, 15]. Weyl’s law describes the distributions of eigenvalues $\upsilon_{n}$ for the Laplace operator $-\nabla^{2}u_{n}=\upsilon_{n}u_{n}$ in a region $\varOmega\subset\mathbb{R}{}^{d}$ with Dirichlet or Neumann boundary conditions. The number of eigenvalues $N_{d}(\upsilon)$ below $\upsilon$ is asymptotically [14, 15]
$N_{d}(\upsilon)=\frac{w_{d}|\varOmega|\upsilon^{d/2}}{(2\pi)^{d}}\quad\text{as}\upsilon\to\infty$ (1)
@@
$N_{d}(k^{2})=\frac{w_{d}|\varOmega|k^{d}}{(2\pi)^{d}}=\frac{w_{d}|\varOmega|}{\lambda^{d}}$ (3)
@@
$N_{1}\approx\frac{k\ell}{\pi}=\frac{2\ell}{\lambda}\quad\text{and }N_{2}\approx\frac{k^{2}A}{4\pi}=\frac{\pi A}{\lambda^{2}}$ (4)
@@
$N\approx\frac{k^{2}A}{2\pi}=\frac{2\pi A}{\lambda^{2}}$ (5)
@@
Figure 1: The number of eigenvalues for Laplace and Helmholtz operators in a region $\varOmega$ is described by Weyl’s law as applied to waveguides with cross sections in $\mathbb{R}{}^{2}$ and one dimensional line elements in (4).
Weyl’s law (5) estimates the number of propagating modes in EM waveguide structures with PEC walls, but it is more challenging to apply Weyl’s law for radiating structures in $\mathbb{R}{}^{3}$. Starting with a spherical region with radius $a$ which is analytically solvable by expanding the radiated field in spherical waves [17, 18]. In contrast to waveguides, there is no clear cut-off between propagating and evanescent waves. There are $N=2L(L+2)$ spherical modes up to order $L$, and modes of order $L\geq ka$ become increasingly reactive [17]. Using the surface area of the sphere together with $L\approx ka$ corresponds to approximately [2]
$N_{\mathrm{A}}\approx 2(ka)^{2}=\frac{k^{2}A}{2\pi}=\frac{2\pi A}{\lambda^{2}}$ (6)
@@
An explanation for the absence of a clear transition between propagating and evanescent modes in the context of a sphere arises from its expanding surface area as the radial distance increases. Consequently, there’s a proportional rise in the number of propagating modes with radial distance. This stands in sharp contrast to waveguiding structures, where the cross-sectional area, and thus the number of degrees of freedom (NDoF), remains constant along the propagation direction. Hence, the approximation $L\approx ka$ emerges as a useful descriptor for the shift from predominantly propagating to evanescent waves.
Although the cutoff $L\approx ka$ provides a good simple approximation for the order of radiating spherical waves. Improved estimates for near-field measurements [19] and computational techniques [20] can be used for more precise estimates. An alternative interpretation of the cutoff is that high-order modes are associated with large current amplitudes and hence low efficiency for radiators of finite conductivity.
Evaluating the number of radiating spherical waves with efficiency above a given threshold, here $50\%$, produces an estimate of the NDoF. Fig. 2 depicts the number of such radiating modes [8] on a spherical shell with surface resistivity $R_{\mathrm{s}}=10^{-p}\eta_{0}$ and spherical ball with resistivity $k\rho_{\mathrm{r}}=10^{-p}\eta_{0}$ for $p=3,5,7$. Note that copper has a surface resistivity of approximately $R_{\mathrm{s}}\approx 0.01\mathrm{\,\Omega/\square}$ around $1\mathrm{\,GHz}$ and a resistivity of $\rho_{\mathrm{r}}\approx 10^{-8}\mathrm{\,\Omega m}$. The results are compared with the estimate (6), and it is observed that the relative difference between the computed NDoF and $2(ka)^{2}$ decreases as $ka\to\infty$. Using material losses opens a possibility to generalize the NDoF to arbitrary shaped radiators as analysed in this paper.
@@
For arbitrary shaped regions with smooth boundary, we can interpret Weyl’s law locally, resulting in a NDoF according to (5) with surface area $A$. However, these DoF will generally only propagate to the volume outside $\varOmega$ and not to the far field. Instead, the NDoF can be considered to propagate from the region $\varOmega$ to a slightly enlarged version of $\varOmega$ as illustrated in Fig. 3a. This short distance is considered fixed and translates to infinity many wavelength in the high-frequency limit. Also note that the NDoF is unbounded if this distance is removed.
Arbitrarily shaped closed surfaces and disjoint regions are much more complex, for which it is also apparent that there are several alternatively ways to interpret DoF, such as communication between regions [3, 21], from a region to a volume surrounding the region in Fig. 3a, and from a region to the far field in Fig. 3b. In this paper, we consider radiating NDoF interpreted as the NDoF from a region to the far field. We also interpret the NDoF from Weyl’s law (3) as the total NDoF that can propagate a short distance from the considered region, as shown in Fig. 3.
@@ III Capacity limits and radiation modes
Treating arbitrary shaped (non-convex or non connected) regions $\varOmega\in\mathbb{R}{}^{3}$, we can formulate a communication problem between transmitters in the region $\varOmega$ to receivers in the far field, or equivalently, on a circumscribing sphere, as shown in Fig. 3b. Current density $\boldsymbol{J}({\boldsymbol{r}})$ in the transmitting region $\varOmega$ constitutes sources for the radiated field. These currents can be expanded in sufficiently large number of basis functions, with expansion coefficients collected in a column matrix $\mathbf{I}$ [22]. Similarly, the radiated field is expanded in spherical waves [17, 18], with expansion coefficients collected in a column matrix $\mathbf{f}=-\mathbf{U}\mathbf{I}$, where $\mathbf{U}$ denotes the projection of spherical waves onto the used basis functions [23]. The capacity (spectral efficiency) of this idealized system $\mathbf{f}=-\mathbf{U}\mathbf{I}+\mathbf{n}$ is determined by [8]
@@
where $\mathbf{P}=\mathcal{E}\{\mathbf{I}\mathbf{I}^{\text{H}}\}$ denotes the covariance matrix of the currents, $\mathbf{R}$ is the resistance matrix [22], and $\gamma$ quantifies the signal-to-noise ratio (SNR) from the additive noise $\mathbf{n}$. In (7), the current is normalized to unit dissipated power [8]. To provide more realistic values for the capacity, physical properties can be considered. For antennas, it is common to incorporate limitations originating in the efficiency [24, 8] or bandwidth from the reactive energy around the antenna [24, 25].
@@
The resistance matrix $\mathbf{R}=\mathbf{R}_{0}+\mathbf{R}_{\rho}$ is here decomposed into the radiation matrix $\mathbf{R}_{0}=\mathbf{U}^{\text{H}}\mathbf{U}$ modelling radiated power $P_{\mathrm{r}}=\frac{1}{2}\mathbf{I}^{\text{H}}\mathbf{R}_{0}\mathbf{I}$ and the material matrix $\mathbf{R}_{\rho}$ modelling Ohmic or dielectric losses $P_{\mathrm{\rho}}=\frac{1}{2}\mathbf{I}^{\text{H}}\mathbf{R}_{\rho}\mathbf{I}$. Radiating modes diagonalizes (8) which solved using waterfilling over radiation mode efficiencies $\nu_{n}$ results in
@@
with a finite number of non-zero power levels $\tilde{P}_{n}$ associated with radiation modes of sufficiently high efficiency. The number of used modes depends on the signal-to-noise ratio $\gamma$ and the distribution of the modes $\nu_{n}$. Waterfilling solutions for five efficiencies $\nu_{n}\in\{1,0.75,0.5,0.1,0.01\}$ and SNRs $\gamma\in\{5,10,20\}$ are depicted in Fig. 4. The allocation of normalized power $\tilde{P}_{n}\gamma$ and number of used modes depend on the SNR $\gamma$. For low $\gamma\to 0$, only the strongest modes are used, and for large $\gamma\to\infty$, all modes are used. For modes with efficiency $\nu_{n}>0.5$, the relative difference between the modes is below $2$, and all these modes can be used for reasonable SNR $\gamma$. For lower efficiencies $\nu_{n}$, $1/\nu_{n}$ increases rapidly, and very large SNR $\gamma$ are needed for the water-filling solution to use these modes, as shown in Fig. 4.
Figure 4: Illustration of water filling solution (10) for radiation mode efficiencies $\nu_{n}\in\{1,0.75,0.5,0.1,0.01\}$.
@@
Geometry and material properties are encapsulated within the radiation modes (9). To facilitate analysis, it is advantageous to distinguishing between the radiation contribution $\mathbf{R}_{0}$ and the material contribution $\mathbf{R}_{\rho}$. Radiation mode eigenvalues are defined by the generalized eigenvalue problem [28]
@@
which can be used to define NDoF from sufficiently efficient modes. Setting a threshold of $50\%$ efficiency corresponding to radiation modes $\varrho_{n}\geq 1$, which is here used to define a NDoF for arbitrary shaped regions and material losses [8].
Figure 5: Normalized radiation modes for spherical shells and spherical balls of the electrical sizes $ka=50$ and $ka=500$.
Fig. 5 depicts normalized radiation modes for spherical shells and spherical balls of the electrical sizes $ka\in\{50,500\}$. The spherical shell has surface resistivity $R_{\mathrm{s}}$, and the spherical ball has resistivity $\rho_{\mathrm{r}}$. Both the surface and volumetric objects exhibit similar normalized radiation modes, each showing a cutoff around $n\approx N_{\mathrm{A}}=2(ka)^{2}$ (6). The cutoff becomes tighter with increasing electrical size, with the $ka=500$ case approaching a step function for the spherical ball.
Alternatively to using a threshold for the efficiency the effective NDoF [29, 30] can be used which expressed in radiation mode efficiency is
@@
The effective NDoF depend on the material losses but not on a threshold level for the efficiency.
Figure 6: Normalized NDoF from radiation modes $\varrho_{n}\geq 1$ and (13) for spherical balls with resistivity $\rho_{\mathrm{r}}=10^{-n}\eta_{0}/k$.
Fig. 6 shows the NDoFs in Fig. 2 normalized with estimate $N_{\mathrm{A}}=2(ka)^{2}$ for the volumetric spherical region. The NDoF for electrically smaller regions are above the $N_{\mathrm{A}}=2(ka)^{2}$ estimate (6) but approaches it for larger sizes. The two definitions based on $\varrho_{n}\geq 1$ and (13) agree well but the $\varrho_{n}\geq 1$ formulation oscillates due to the onset of modes at specific frequencies.
Radiation modes and hence NDoF defined by (11) or (13) can be calculated for arbitrary shapes and material parameters. Fig. 7 shows an example of NDoF for the six different shapes in Tab. I modelled by a surface resistivity $R_{\mathrm{s}}=10^{-5}\eta_{0}$. It is observed that the NDoF increase with the electrical size $ka$, with the spherical shell exhibiting the highest NDoF, followed by the solid cylinder and the corrugated cylinder, which have slightly higher NDoF than the open cylinder and connected discs. The single disc exhibits the lowest NDoF. The effective NDoF (13) produces similar results but with less oscillations. This paper demonstrates that the asymptotic number of NDoF is proportional to the electrical size of the region’s average shadow area $\langle A_{\mathrm{s}}\rangle$ in Tab. I, see also [10] for the corresponding number of significant characteristic modes.
Figure 7: NDoF from radiation modes $\varrho_{n}\geq 1$ in (11) and $N_{\mathrm{e}}$ in (13) for the six objects in Tab. I with surface resistivity $R_{\mathrm{s}}=10^{-5}\eta_{0}$.
TABLE I: Average shadow area $\langle A_{\mathrm{s}}\rangle$, surface area $A$, and height to radius ratio for a sphere, cylinder, disc, open cylinder, corrugated cylinder, and connected discs. Objects are circumscribed by a sphere with radius $a$.

### v2

@@ Abstract
Electromagnetic degrees of freedom are instrumental in antenna design, wireless communications, imaging, and scattering. Larger number of degrees of freedom enhances control in antenna design, influencing radiation patterns and directivity, while in communication systems, it links to spatial channels for increased data rates and reliability, and resolution in imaging. The correlation between computed degrees of freedom and physical quantities is not fully understood, prompting a comparison between classical estimates, Weyl’s law, modal expansions, and optimization techniques. In this paper, it is shown that the number of degrees of freedom for arbitrary shaped radiating structures approaches the shadow area measured in squared wavelengths asymptotically as the wavelength decreases.
@@ I Introduction
Electromagnetic degrees of freedom (DoF) finds importance across a wide range of electromagnetic (EM) applications, spanning antenna design, wireless communication systems, scattering problems, and measurement techniques [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. The number of DoF (NDoF) emerges as a crucial factor in achieving desired performance levels. For instance, in antenna design, a higher NDoF facilitates enhanced control over radiation patterns and maximum directivity [6]. In wireless communication systems, NDoF correlates with the number of channels or spatial dimensions utilized for signal transmission, thereby providing higher data rates and improved reliability, especially in technologies like multiple-input multiple-output (MIMO) systems and intelligent surfaces [9, 7].
@@
The relationship between the NDoF and the electrical size of an object have been explored for at least half-a-century investigating resolution in imaging systems [12] and antenna systems [13]. Small electrical sizes exhibit few DoF while large sizes display numerous DoF and hence, e.g., potentially higher directivity and narrow beamwidth. Analytically solvable cases such as spherical regions with radius $a$ offer a simple estimate of $2(ka)^{2}$ DoFs [2] based on the first $L\approx ka$ spherical modes for wavenumber $k=2\pi/\lambda$, wavelength $\lambda$, and two polarization DoFs. The number of propagating waves in hollow rectangular waveguides [14] are also solvable with asymptotically $2\pi A/\lambda^{2}$ DoFs for a cross-sectional area $A$. These results are traced back to Weyl more than a century ago, who studied the eigenvalue distribution of the Laplace and Helmholtz operators [15, 16]. Alternative estimates based on Nyquist ($\lambda/2$) sampling resulting in $8A/\lambda^{2}$ DoFs have also been extensively used [2, 17, 4, 5, 18]. Moreover, communication systems have recently been investigated for different configurations [9, 7], with $2\pi A/\lambda^{2}$ DoFs for planar apertures.
In this paper, Weyl’s law [15, 16] and its connection to the NDoF are first explored. Beyond its mathematical significance in spectral theory, Weyl’s law is also fundamental in explaining the NDoF in thermal radiation [16]. However, its relevance to the NDoF in antenna and communication systems has been largely overlooked. Weyl’s law demonstrates that the NDoF for arbitrarily shaped bodies scale with the surface area, measured in squared wavelengths. These NDoF can be interpreted as the DoF in a communication system between the body and the surrounding region. Weyl’s law provides a unified framework that extends previous results on the asymptotic NDoF of canonical shapes, such as spheres [2] and planar rectangles [9, 7], to arbitrarily shaped regions.
In antenna systems, it is more common to focus on radiating systems, where the antenna transmits energy to the far-field. This paper shows that the radiating NDoF approaches the total shadow area of the region, measured in squared wavelengths. For convex shapes, this result matches the NDoF derived from Weyl’s law, but it differs for more general shapes. The reduction in NDoF for non-convex shapes can be attributed to the blockage of some regions, which limits their contribution to the radiated fields.
The NDoF for radiating systems is crucial in antenna and communication system design, as it determines the maximum number of radiating fields that can be achieved within a constrained design region. Furthermore, it is directly related to the maximum achievable average gain of an antenna system when its beam is steered across all possible directions. This establishes a connection to the Hannan limit, which was originally derived for infinite planar arrays [19, 20]. Notably, this radiating NDoF is also twice the number of significant characteristic modes [10]. In communication systems, the derived asymptotic limit offers a straightforward estimate of the NDoF, and thus the capacity [18]. This generalizes the results previously obtained for planar rectangular regions [9, 7] to arbitrarily shaped regions. The possibility of extending this framework to communication between arbitrarily shaped regions has also been explored [21].
The paper is organized as follows. Section II introduces Weyl’s law and its applications for determining NDoF. Section III explores the concept of capacity and its connection to radiation modes and NDoF. In Section IV, the asymptotic NDoF is analyzed in relation to the shadow area of a region. Electric and magnetic currents are examined in Section V. NDoF in inverse problems is discussed in Section VI. Finally, the paper is concluded in Section VII.
@@ II Weyl’s law and propagating modes
NDoF is most straight forward determined for the number of propagating modes in a waveguide. Canonical shapes such as rectangular or circular cross sections can be solved analytically [14]. These and arbitrary shaped cross sections can also be treated as a special case of Weyl’s law [15, 16]. Weyl’s law describes the distributions of eigenvalues $\nu_{n}$ for the Laplace operator $-\nabla^{2}u_{n}=\nu_{n}u_{n}$ in a region $\varOmega\subset\mathbb{R}{}^{d}$ with Dirichlet or Neumann boundary conditions. The number of eigenvalues $N_{\mathrm{W}d}(\nu)$ below $\nu$ is asymptotically [15, 16]
$N_{\mathrm{W}d}(\nu)=\frac{w_{d}|\varOmega|\nu^{d/2}}{(2\pi)^{d}}\quad\text{as}\nu\to\infty$ (1)
@@
$N_{\mathrm{W}d}(k^{2})=\frac{w_{d}|\varOmega|k^{d}}{(2\pi)^{d}}=\frac{w_{d}|\varOmega|}{\lambda^{d}}$ (3)
@@
$N_{\mathrm{W}1}=\frac{k\ell}{\pi}=\frac{2\ell}{\lambda}\quad\text{and }N_{\mathrm{W}2}=\frac{k^{2}A}{4\pi}=\frac{\pi A}{\lambda^{2}}$ (4)
@@
$N_{\mathrm{A}}=\frac{k^{2}A}{2\pi}=\frac{2\pi A}{\lambda^{2}}$ (5)
@@
Figure 1: The number of eigenvalues for Laplace and Helmholtz operators in a region $\varOmega$ is described by Weyl’s law as applied to waveguides and line elements in (4).
Weyl’s law (5) estimates the number of propagating modes in electromagnetic waveguide structures with perfect electric conductor (PEC) walls, but it is more challenging to apply Weyl’s law for radiating structures in $\mathbb{R}{}^{3}$. Starting with a spherical region with radius $a$ which is analytically solvable by expanding the radiated field in spherical waves [22, 23]. In contrast to waveguides, there is no clear cut-off between propagating and evanescent waves. There are $2L(L+2)$ spherical modes up to order $L$, and modes of order $L\geq ka$ become increasingly reactive [22]. Using the surface area of the sphere together with $L\approx ka$ corresponds to approximately [2]
$2(ka)^{2}=\frac{k^{2}A}{2\pi}=\frac{2\pi A}{\lambda^{2}}$ (6)
@@
An explanation for the absence of a clear transition between propagating and evanescent modes in the context of a sphere arises from its expanding surface area as the radial distance increases. Consequently, there is a proportional rise in the number of propagating modes with radial distance. This stands in sharp contrast to waveguiding structures, where the cross-sectional area, and thus the NDoF, remains constant along the propagation direction. Hence, the approximation $L\approx ka$ emerges as a useful descriptor for the shift from predominantly propagating to evanescent waves.
Although the cutoff $L\approx ka$ provides a good simple approximation for the order of radiating spherical waves. Improved estimates for near-field measurements [24] and computational techniques [25] can be used for more precise estimates. An alternative interpretation of the cutoff is that high-order modes are associated with large current amplitudes and hence low efficiency for radiators of finite conductivity. This is related to the rapidly increasing reactive near field for high order modes, i.e., there are large field (and current) amplitudes that do not contribute efficiently to the radiated power. These fields are associated with high losses and stored energy for antennas. This imbalance between the amplitude of the modes can also be interpreted as a communication channel with large dynamical range between the singular values of the channel matrix requiring a high SNR.
@@
Evaluating the number of radiating spherical waves with efficiency above a given threshold, here $50\%$, produces an estimate of the NDoF. Fig. 2 depicts the number of such radiating modes [8] on a spherical shell with surface resistivity $R_{\mathrm{s}}=10^{-p}\eta_{0}$ and spherical ball with resistivity $k\rho_{\mathrm{r}}=10^{-p}\eta_{0}$ for $p=3,5,7$. Note that copper has a surface resistivity of approximately $R_{\mathrm{s}}\approx 0.01\mathrm{\,\Omega/\square}$ around $1\mathrm{\,GHz}$ and a resistivity of $\rho_{\mathrm{r}}\approx 10^{-8}\mathrm{\,\Omega\,m}$. The results are compared with the estimate (6), and it is observed that the relative difference between the computed NDoF and $2(ka)^{2}$ decreases as $ka\to\infty$. The radiation efficiencies for the spherical regions used in Fig. 2 are determined by expanding the current density and fields in spherical waves. The solution is expressed in radiation modes [26, 27] as given in App. A. Using material losses opens a possibility to generalize the NDoF to arbitrary shaped radiators as analyzed in this paper.
Arbitrarily shaped closed surfaces and disjoint regions are much more complex, for which it is also apparent that there are several alternative ways to interpret DoF, such as communication between regions [3, 28, 11, 21], from a region to a volume surrounding the region in Fig. 3a, and from a region to the far field in Fig. 3b. In this paper, we consider radiating NDoF interpreted as the NDoF from a region to the far field.
For arbitrary shaped regions with smooth boundary, we can interpret Weyl’s law locally, resulting in a NDoF according to (5) with surface area $A$. However, these DoF will generally only propagate to the volume outside $\varOmega$ and not to the far field. Instead, the NDoF can be considered to propagate from the region $\varOmega$ to a slightly enlarged version of $\varOmega$ as illustrated in Fig. 3a. This short distance is considered fixed and translates to infinity many wavelengths in the high-frequency limit. Also note that the NDoF is unbounded if this distance is removed.
@@ III Capacity limits and radiation modes
Treating arbitrary shaped (non-convex or non-connected) regions $\varOmega\in\mathbb{R}{}^{3}$, we can formulate a communication problem between transmitters in the region $\varOmega$ to receivers in the far field, or equivalently, on a circumscribing sphere, as shown in Fig. 3b. Current density $\boldsymbol{J}({\boldsymbol{r}})$ in the transmitting region $\varOmega$ constitutes sources for the radiated field. These currents can be expanded in sufficiently large number of basis functions, with expansion coefficients collected in a column matrix $\mathbf{I}$ [29]. Similarly, the radiated field is expanded in spherical waves [22, 23], with expansion coefficients collected in a column matrix $\mathbf{f}=-\mathbf{U}\mathbf{I}$, where $\mathbf{U}$ denotes the projection of spherical waves onto the used basis functions [30]. The capacity (spectral efficiency) of this idealized system $\mathbf{f}=-\mathbf{U}\mathbf{I}+\mathbf{n}$ is determined by [8]
@@
where $\mathbf{P}=\mathcal{E}\{\mathbf{I}\mathbf{I}^{\text{H}}\}$ denotes the covariance matrix of the currents, $\mathbf{R}$ is the resistance matrix [29], and $\gamma$ quantifies the signal-to-noise ratio (SNR) from the additive noise $\mathbf{n}$. In (7), the current is normalized to (two times) unit dissipated power [8]. To provide more realistic values for the capacity, physical properties can be considered. For antennas, it is common to incorporate limitations originating in the efficiency [31, 8] or bandwidth from the reactive energy around the antenna [31, 32].
@@
The resistance matrix $\mathbf{R}=\mathbf{R}_{0}+\mathbf{R}_{\rho}$ is here decomposed into the radiation matrix $\mathbf{R}_{0}=\mathbf{U}^{\text{H}}\mathbf{U}$ modelling radiated power $P_{\mathrm{r}}=\frac{1}{2}\mathbf{I}^{\text{H}}\mathbf{R}_{0}\mathbf{I}$ and the material matrix $\mathbf{R}_{\rho}$ modelling Ohmic or dielectric losses $P_{\mathrm{\rho}}=\frac{1}{2}\mathbf{I}^{\text{H}}\mathbf{R}_{\rho}\mathbf{I}$.
Figure 4: Illustration of the water filling solution (10) for radiation mode efficiencies $\nu_{n}\in\{1,0.75,0.5,0.1,0.01\}$. The red bars represent $1/\nu_{n}$, while the hashed blue bars indicate the allocated power $\tilde{P}_{n}\gamma$ for the different modes.
Radiating modes diagonalizes (8), which solved using waterfilling over radiation mode efficiencies $\nu_{n}$ results in
@@
with a finite number of non-zero power levels $\tilde{P}_{n}\geq 0$ associated with radiation modes of sufficiently high efficiency. The number of used modes depends on the signal-to-noise ratio $\gamma$ and the distribution of the modes $\nu_{n}$. Waterfilling solutions for five efficiencies $\nu_{n}\in\{1,0.75,0.5,0.1,0.01\}$ and SNRs $\gamma\in\{5,10,20\}$ are depicted in Fig. 4. The reciprocals of the efficiencies, $1/\nu_{n}$, are large for inefficient modes, e.g., mode 4 has $101/\nu_{4}=$ 10, and mode 5 has $1001/\nu_{5}=$ 100. The normalized power level allocated to each mode, $\tilde{P}_{n}\gamma$, in the water-filling algorithm can be interpreted as determining the water level over a bottom profile defined by $1/\nu_{n}$, with a total amount of water $\gamma=\sum\tilde{P}_{n}\gamma$. The allocation of normalized power $\tilde{P}_{n}\gamma$ and the number of active modes depend on the SNR, $\gamma$. For low SNR, $\gamma\ll 1$, only the strongest mode, with $11/\nu_{1}=$ 1, contributes. As $\gamma$ increases, power is quickly allocated to additional modes, such as modes 2 and 3, with $1/\nu_{2}\approx 1.33$ and $21/\nu_{3}=$ 2, respectively. Inefficient modes, such as $101/\nu_{4}=$ 10, require significantly higher SNR to be activated, as shown in Fig. 4.
@@
Geometry and material properties are encapsulated within the radiation modes (9). To facilitate analysis, it is advantageous to distinguish between the radiation contribution $\mathbf{R}_{0}$ and the material contribution $\mathbf{R}_{\rho}$. Radiation mode eigenvalues are defined by the generalized eigenvalue problem [26]
@@
which can be used to define NDoF from sufficiently efficient modes. Setting a threshold of $50\%$ efficiency corresponding to radiation modes $\varrho_{n}\geq 1$, which is here used to define a NDoF $N_{1}$ for arbitrary shaped regions and material losses [8].
Figure 5: Normalized radiation modes for spherical shells and spherical balls of the electrical sizes $ka=50$ and $ka=500$. The mode index is normalized by the asymptotic NDoF $N_{\mathrm{A}}=2(ka)^{2}$.
Figure 6: Comparison between the NDoF $N_{1}$ defined by $50\%$ efficiency ($\varrho_{n}\geq 1$) and the effective NDoF $N_{\mathrm{e}}$ in (13). The NDoFs are normalized by $N_{\mathrm{A}}=2(ka)^{2}$ for spherical balls with resistivity $\rho_{\mathrm{r}}=10^{-n}\eta_{0}/k$ with $n=3,5,7$.
Fig. 5 depicts normalized radiation modes for spherical shells and spherical balls of the electrical sizes $ka\in\{50,500\}$, see App. A. The spherical shell has surface resistivity $R_{\mathrm{s}}$, and the spherical ball has resistivity $\rho_{\mathrm{r}}$. Both the surface and volumetric objects exhibit similar normalized radiation modes, each showing a cutoff around $n\approx N_{\mathrm{A}}=2(ka)^{2}$. The cutoff becomes tighter with increasing electrical size, with the $ka=500$ case approaching a step function for the spherical ball.
Alternatively, instead of using a threshold for efficiency, the effective NDoF [34, 35] can be utilized, which, when expressed in terms of radiation mode efficiency, is given by
@@
Fig. 6 compares the NDoF $N_{1}$ based on $50\%$ efficiency, as in Fig. 2, with the asymptotic effective NDoF $N_{\mathrm{e}}$ in (13). The depicted NDoFs are normalized with the estimate $N_{\mathrm{A}}=2(ka)^{2}$ for the spherical region. The NDoF based on $\varrho_{n}\geq 1$ and the NDoF  (13) agree well but the $\varrho_{n}\geq 1$ formulation oscillates due to the onset of modes at specific frequencies. The NDoFs for electrically smaller regions are above the $N_{\mathrm{A}}=2(ka)^{2}$ estimate (6) but approaches it for larger sizes.
Figure 7: NDoF $N_{1}$ from radiation modes $\varrho_{n}\geq 1$ in (11) and $N_{\mathrm{e}}$ in (13) for the six objects in Tab. I with surface resistivity $R_{\mathrm{s}}=10^{-5}\eta_{0}$.
Radiation modes and hence NDoF defined by (11) or (13) can be calculated for arbitrary shapes and material parameters. Fig. 7 shows an example of NDoF for the six different shapes in Tab. I modelled by a surface resistivity $R_{\mathrm{s}}=10^{-5}\eta_{0}$. It is observed that the NDoF increase with the electrical size $ka$, with the spherical shell exhibiting the highest NDoF, followed by the solid cylinder and the corrugated cylinder, which have slightly higher NDoF than the open cylinder and connected discs. The single disc exhibits the lowest NDoF. The effective NDoF (13) produces similar results but with fewer oscillations. This paper demonstrates that the asymptotic number of NDoF is proportional to the electrical size of the region’s average shadow area $\langle A_{\mathrm{s}}\rangle$ in Tab. I, see also [10] for the corresponding number of significant characteristic modes.
TABLE I: Average shadow area $\langle A_{\mathrm{s}}\rangle$, surface area $A$, and height to radius ratio $h/r$ for a sphere, cylinder, disc, open cylinder, corrugated cylinder, and connected discs. Objects are circumscribed by a sphere with radius $a$.